OpenStudy (anonymous):

Fill in the value of x. log^81(1/9)=x

hartnn (hartnn):

is it \(\huge \log_{81}(1/9)=x ??\)

hartnn (hartnn):

then u just need 2 properties of log \(\large log_ab=\frac{log b}{log a}\\ \large log a^n=nloga\)

OpenStudy (anonymous):

i think i fpund new properties \[\log_{a^b }(c) = \frac{1}{b}\log_{a}(c) \] @hartnn can you verify

hartnn (hartnn):

that is correct ! good work :)

hartnn (hartnn):

could u solve further @smourin ?